4 A car accelerates uniformly in a straight line from 10 m/s to 20 m/s in 5 seconds. What is the acceleration?

4

A car accelerates uniformly in a straight line from 10 m/s to 20 m/s in 5 seconds. What is the acceleration?

Grade:12

1 Answers

Nirmal Singh.
10 years ago
we know that v = u + at
20 = 10 +a *5
5 a = 20-10
5a = 10
a =2 m/s^2
Thanks & Regards,
Nirmal Singh
Askiitians faculty